OpenStudy (anonymous):

I need to know how to work: 2(x+4)^2=9

OpenStudy (anonymous):

hm

OpenStudy (anonymous):

2(x+4)²=9, this is as similar as 2x²+16x+32=9, so, 2x²+16x +23=0. Now use Bhaskara´s equation and find the x´s values